Circular No. 9111

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

U SCORPII B. E. Schaefer, Louisiana State University, reports that B. G. Harris

(New Smyrna Beach, FL, USA) has discovered the recurrent nova U Sco (cf. IAUC 7749, 8279) to be in outburst. On Jan. 28.4385 UT, U Sco was recorded with CCD images and visual observations to be at magnitude V = 8.05. This discovery has been confirmed by B. E. Schaefer with a 15.2-cm telescope and by S. Dvorak (Clermont, FL, USA) at mag ∼ 8.8 visually. M. Templeton, AAVSO, reports the following previous magnitudes for U Sco: Jan. 27.450, V = 18.2 (Harris); 27.627, [16.5 (M. Linnolt, Hoolehua, HI, USA; visual). This is the tenth discovered eruption, with the last being in 1999 (cf. IAUC 7113), when it reached visual mag 7.5. U Sco is located at α = 16^{h}22^{m}30^{s}.80, δ = −17^{o}52′43′′.0 (equinox 2000.0). The entire peak (from quiescence to peak, to one mag below peak) occurring inside of one day. This fast event requires fast reaction to get observations near the peak.

V2673 OPHIUCHI = NOVA OPHIUCHI 2010 As first reported on CBETs 2128 and 2139 (where additional details

may be found), S. Nakano, Sumoto, Japan, forwarded the discovery by H. Nishimura (Miyawaki, Kakegawa, Shizuoka-ken, Japan) of an apparent nova (mag 8.4) on frames taken on Jan. 15.857 UT with a Canon EOS 5D camera (+ Minolta 120-mm f/3.5 lens); Nakano measured mag 8.8 and the following position from the discovery image: α = 17^{h}39^{m}40^{s}.90, δ = −21^{o}39′50′′.5 (equinox 2000.0). Additional CCD magnitudes: 2009 Nov. 3, [11.0 (Nishimura); 2010 Jan. 13.86, [9.5: (Nishimura); 14.865, 10.1 (Nishimura; measured by Nakano; limiting mag 10.6); 16.856, 8.4 (K. Itagaki, Takanezawa station, Tochigi-ken, Japan, 0.30-m reflector; position end figures 40^{s}.94, 47′′.9); 16.860, 8.2 (K. Kadota, Ageo, Saitama-ken, Japan, 0.25-m reflector; position end figures 40^{s}.97, 47′′.4); 18.299, 8.4 (W. Vollmann, Vienna, Austria, and H. Koberger, Vilaflor, Tenerife; green-band); 22.250, 9.5 (Vollmann and Koberger). Kadota adds that nothing is visible at this position on a red Digitized Sky Survey image from 1997 (limiting mag estimated to be 20 by Nakano). H. Maehara, Kwasan Observatory, Kyoto University, reports that a low-resolution spectrum (range 400–800 nm), obtained on Jan. 22.890 with a 25-cm telescope at Kwasan Observatory, shows Hα and Fe II emission lines and suggests that this object is a classical nova just after maximum. E. Kazarovets and N. Samus report that the GCVS designation V2673 Oph has been assigned to this nova.

2010 January 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green